Potassium permanganate(KMNO4)is has a solubility of 6.4 g/ 100 g of water at 20ºC, and 250 g / Literof water at 65ºC(Use 1.00 g/mL as the density of water)

a)Can you dissolve more or less potassium permanganateby increasing the temperature of the solution? Clearly show your calculations or reasoning.

b)What mass of potassium permanganatecan dissolve in 65 g of water at 20ºC?

c)What mass of potassium permanganate will precipitate out of a solution made by dissolving 21.3 g of potassium permanganate in 65 g of water at 65ºC and then cooling the solution down to 20ºC?
Answer #1

b)

At 20 deg C,

Mass of KMnO4 dissolve in 100 g water = 6.4 g

Mass of KMnO4 dissolve in 65 g water = ( 6.4 x 65 ) / 100

Mass of KMnO4 dissolve in 65 g water = 4.16 g

c)

At 65 deg C,

Mass of KMnO4 dissolve in 1000 g water = 250 g

Mass of KMnO4 dissolve in 65 g water = ( 250 x 65 ) / 1000

Mass of KMnO4 dissolve in 65 g water = 16.25 g

Mass of KMnO4 left = 21.3 - 16.25 = 5.05 g

Mass of KMnO4 dissolve in 65 g water at 20 deg C = 4.16 g

Mass of KMnO4 left at 20 deg C = 5.05 - 4.16 = 0.89 g